7. Mr. M. is admitted to the hospital with a GI bleed and his physician orders a blood transfusion of packed red blood cells (PRBCs). Order: 1 unit PRBCs; 520 ml IV to infuse over 4 hours. Drop factor: 20 gtt/ml. What is the infusion rate in gtt/min?
Answer #1

ANSWER.

7) Calculation of infusion rate(gtt/min).

       Formula

           Infusion rate(gtt/min) = Volume(ml)/Time(min) * Drop factor(gtt/ml).

Here

      Volume(ml) = 520 ml.

      Time(min) = 240 min ( 4hrs = 4 * 60 min)

      Drop factor = 20 gtt/ml

As per above formula

        Infusion rate(gtt/min) = 520/240 * 20

                                       = 2.17 * 20

                                         = 43.33 gtt/min.